1. [(7x8) - (18/3) + (-2) x (-4)] x (6/3)
2. (1/2 +1/4) divide by 4/9
1
Expert's answer
2020-09-10T19:21:31-0400

1. "[(7\times8)-(" "18\over3" ")" "+" "(-2)\times(-4)" "]" "\times" "(" "6\over 3" ")"

"=(56-6-2\times 4) \times 2"

"=(56-6+8) \times 2"

"=(50+8)\times 2"

"58\times 2 =116"


2."(" "1\over 2" "+" "1\over 4" ")" "\div" "4\over9"

"1\over 2" "+" "1\over 4" "=" "3\over4"

"3\over 4" "\times" "9\over 4" "=" "27\over 16"


"=1" "11\over 16"

.